OpenStudy (anonymous):

If the simplified expression is written in standard form, what is the leading coefficient? x√8x^3+x√2+4√2x

OpenStudy (campbell_st):

well the 1st term can be rewritten \[x \sqrt{8x^3} = x \sqrt{4x^2 \times 2x}=x \times 2x \sqrt{2x} = 2x^2\sqrt{2x}\] so its \[2x^2\sqrt{2x} + 4\sqrt{2x} + x \sqrt{2} = (2x^2 + 4)\sqrt{2x} + x \sqrt{2}\]

OpenStudy (anonymous):

short to say is tht 2 is the leading coefficient